# Monomial web basis for the SL(N) skein algebra of the twice punctured sphere

### v1

@@ 4.4.5. Quantum trace map: definition, part 2 of 2
for $(a,b,c)$ interior; while for boundary vertices:
$P(j(n-j)0,(j+1)(n-j-1)0)$ $=P((n-j)0j,(n-j-1)0(j+1)$
$=P(0j(n-j),0(j+1)(n-j-1))=+1$
@@
For $j=1,2,\dots,n-1$ and symbols $Z$ and $X$, define $n\times n$ diagonal and block diagonal matrices
$S^{\mathrm{edge}}_{j}(Z)$ $=Z^{-j/n}\mathrm{diag}\left(\begin{matrix}Z&Z&\ldots&Z&1&1&\ldots&1\end{matrix}\right),$
$S^{\mathrm{left}}_{j}(X)$ $=X^{-(j-1)/n}\mathrm{diag}\left(\begin{matrix}X&\ldots&X&\left(\begin{matrix}1&1\\ 0&1\end{matrix}\right)&1&\ldots&1\end{matrix}\right),$
$S^{\mathrm{right}}_{j}(X)$ $=X^{+(j-1)/n}\mathrm{diag}\left(\begin{matrix}1&\ldots&1&\left(\begin{matrix}1&0\\ 1&1\end{matrix}\right)&X^{-1}&\ldots&X^{-1}\end{matrix}\right)$
@@
$L^{\omega}_{i,j}=\left[M^{\mathrm{edge}}(Z_{1},\dots,Z_{n-1})M^{\mathrm{left}}((X_{abc})_{abc})M^{\mathrm{edge}}(Z^{\prime}_{1},\dots,Z^{\prime}_{n-1})\right]_{\overline{i},\overline{j}}$
and
@@ 4.5.2. Annulus: a piecewise linear computation
Let $L^{\omega}$ and $R^{\omega}$ be the standard quantum left and right matrices. In general, an entry $L^{\omega}_{ij}$ or $R^{\omega}_{ij}$ is a sum of monomials in $\mathcal{T}_{n}^{\omega}(\mathfrak{D}_{3})$. Put $M=L^{\omega}$, which recall is lower triangular. Note that similar facts will hold for the right matrix as well, but we will not need them.
@@
    $\mathrm{deg}\left(\prod_{i=n+1-k}^{n}M_{ii}\right)(a,b,c)=\min\{f_{k}(n-a),g_{k}(n-a)\}$
    and
@@
  1. (1)
    By definition, $M$ is the product of lower triangular, and diagonal, matrices $M_{v}$, $v\in\Theta_{n}^{\prime}$, each of whose $ij$-entry is of the form $X_{v}^{d_{vij}/n}$ for some $d_{vij}\in\mathbb{Z}$, as well as lower triangular matrices involving no variables. Rather, $M$ is the Weyl ordering of such a product, but the Weyl ordering will play no role throughout this proof since we are only interested in degrees, so we will omit it for notational simplicity.
  2. (2)
    This is one of the relations appearing in the main result of [13].
  3. (3)
    The first statement is immediate from the form of the matrices $M_{v}$. The second statement is because the off diagonal terms $X_{v}^{d_{vij}/n}$ of the $M_{v}$ have exponent $d_{vij}<0$.
  4. (4)
    It is clear that $M$ does not involve the variables $X_{0j(n-j)}$, so assume $a\geq 1$. Recall that the $(\overline{m},\overline{m})$ entry of $M$ is the $(m,m)$ entry of the product of matrices $S^{\mathrm{edge}}_{l}(X_{l0(n-l)})$ and $S^{\mathrm{edge}}_{l}(X_{l(n-l)0}$ for $11\leq l\leq n-$ 1 and $S^{\mathrm{left}}_{j}(X_{(j-1)(n-i)(i-j+1)})$ for $12\leq j\leq i\leq n-$ 1, and $S^{\mathrm{left}}_{1}$ a number of times. We only need to observe then that, in the former, $X_{l0(n-l)}$ and $X_{l(n-l)0}$ appear with power $1-\ell/n$ (resp. $-\ell/n$) when $m\leq l$ (resp. $m>l$), and, in the latter, $X_{(j-1)(n-i)(i-j+1)}$ appears with power $1-(j-1)/n$ (resp. $-(j-1)/n$) when $m\leq j-1$ (resp. $m>j-1$).
  5. (5)
    This is part of the mathematical content of [47, Equation (170) and Theorem 10.2]. For the sake of self-containedness, we provide this short calculation here as well. The formulas work when $a=0$, so assume $a\geq 1$. Note that the larger degree $n-a$ for $M_{ii}$ occurs if and only if $n+1-a\leq i$, else it has the smaller degree $-a$.
@@ 4.5.3. Annulus: finishing
Abusing notation and putting $t^{R}_{k}:\mathbb{R}\to\mathbb{R}$ defined by $t^{R}_{k}(a)=\min\{f_{k}(n-a),g_{k}(n-a)\}$, it suffices to show that the $n-1$ vectors $(t^{R}_{k}(n-1),\dots,t^{R}_{k}(2),t^{R}_{k}(1))$ in $\mathbb{R}^{n-1}$ varying over $k=1,2,\dots,n-1$ are linearly independent. Indeed, if we arrange these $n-1$ vectors as the columns of a $(n-1)\times(n-1)$ matrix $T^{R}$, $T^{R}_{ij}=t^{R}_{j}(n-i)$, then this matrix is the inverse of the matrix $\left(T^{R}\right)^{-1}=(1/n)\left(\begin{matrix}&&&&0&0&-1&2\\ &&&&0&-1&2&-1\\ &&&&-1&2&-1&0\\ \vdots&\vdots&\vdots&\reflectbox{$\ddots$}&&&&\\ -1&2&-1&\ldots&&&&\\ 2&-1&0&\ldots&&&&\end{matrix}\right)$. ∎
@@ 5.1. Embedding the commutative polynomial algebra into quantum tori
In other words, for every such $q^{1/n}$ we obtain an explicit geometric embedding of the commutative polynomial algebra $\mathcal{S}_{n}^{q}\cong\mathbb{C}[x_{1},x_{2},\dots,x_{n-1}]$ into the quantum torus $\mathcal{T}_{n}^{\omega}(\mathfrak{D}_{3}^{L})\otimes\mathcal{T}_{n}^{\omega}(\mathfrak{D}_{3}^{R})$.
@@
By [66, Corollary 20], see also [6, 59] for the case $n=2$, the skein module $\mathcal{S}_{n}^{q}(\mathfrak{M})$ of a three-dimensional manifold $\mathfrak{M}$ (Remark 1.5) evaluated at $q^{1/n}=1$ is isomorphic to the, a priori, unreduced coordinate ring over $\mathbb{C}$ of the $\mathrm{SL}(n)$ character variety of $\mathfrak{M}$. Said another way, the reduced coordinate ring of the $\mathrm{SL}(n)$ character variety of $\mathfrak{M}$ is isomorphic to $\mathcal{S}_{n}^{q}(\mathfrak{M})(q^{1/n}=1)/\sqrt{0}$, namely the quotient of $\mathcal{S}_{n}^{q}(\mathfrak{M})(q^{1/n}=1)$ by its set of nilpotent elements. By Corollary 5.3, since $q^{1/n}=1\notin Q$, taking $\mathfrak{M}=\mathfrak{A}\times\mathfrak{I}$, the coordinate ring of the $\mathrm{SL}(n)$ character variety of the annulus is reduced and isomorphic to the skein algebra $\mathcal{S}_{n}^{q}(q^{1/n}=1)\cong\mathbb{C}[x_{1},x_{2},\dots,x_{n-1}]$.
@@ 5.2. Injectivity of the splitting map for the twice punctured sphere
The argument is identical to that of the $n=3$ result [34, Theorem 8.1], or the earlier $n=2$ result of [45], both proved for any surface $\mathfrak{S}$, once it is known that the monomial webs $(B^{\circ}_{1})^{\prime m_{1}}(B^{\circ}_{2})^{\prime m_{2}}\cdots(B^{\circ}_{n-1})^{\prime m_{n-1}}$ in the stated skein algebra $\widetilde{\mathcal{S}}_{n}^{q}(\overline{\mathfrak{S}})$, which are the images under the splitting map of the monomial basis webs $(B^{\circ}_{1})^{m_{1}}(B^{\circ}_{2})^{m_{2}}\cdots(B^{\circ}_{n-1})^{m_{n-1}}$ in $\left(\mathcal{S}_{n}^{q}\right)^{\circ}$, are linearly independent in $\widetilde{\mathcal{S}}_{n}^{q}(\overline{\mathfrak{S}})$. This was essentially proved above as part of the proof of the main theorem. Technically, this independence was proved for the annular surface $\overline{\overline{\mathfrak{S}}}$ obtained from $\overline{\mathfrak{S}}$ by cutting out another monoangle from the remaining internal puncture, but since the monomial webs in the surface $\overline{\overline{\mathfrak{S}}}$ are the images under the splitting map of the monomial webs in the surface $\overline{\mathfrak{S}}$, the latter are independent as well. Note also that, when mimicking the argument of [34], the $n=3$ ‘cut’ and ‘glue’ maps of [34, Section 7] are replaced by the corresponding $n=n$ maps in [46, Section 7.4]. ∎
@@ 6. Fock–Goncharov duality
The balanced Knutson–Tao lattice (resp. fan) is the set of functions $f:\Theta_{n}^{\prime}\to\mathbb{R}$ satisfying the property that all the rhombus numbers $r(f)^{\mathrm{t}}_{ij}$ and $r(f)^{\mathrm{bl}}_{ij}$ and $r(f)^{\mathrm{br}}_{ij}$ are integers (resp. natural numbers). Note that the balanced Knutson–Tao lattice (resp. fan) is a subgroup (resp. sub-monoid) of $\mathbb{R}^{\Theta_{n}^{\prime}}$ under addition, in particular it contains the zero function. We denote by $\mathcal{C}$ the sub-monoid that is the intersection of the balanced Knutson–Tao fan with those functions $f$ satisfying $f(a,b,c)=0$ for all $a=0$.
Let the functions $t^{L}_{k}$ and $t^{R}_{k}$ in $\mathbb{N}^{\Theta_{n}^{\prime}}$ for $k=1,2,\dots,n-1$ be defined as in Section 4.5.3. One checks that $t^{R}_{k}\in\mathcal{C}$ for each $k$. By the symmetry $t^{L}_{n-k}=t^{R}_{k}$ this is also true for the $t^{L}_{k}$. Moreover, we have the following specific values for the rhombus numbers: $r(t^{R}_{k})^{\mathrm{bl}}_{kj}=1$ for all $j=1,2,\dots,n-k$ and all other rhombus numbers $r(t^{R}_{k})^{\mathrm{bl}}_{ij}$, $i\neq k$, and $r(t^{R}_{k})^{\mathrm{t}}_{ij}$ and $r(t^{R}_{k})^{\mathrm{br}}_{ij}$ are zero.
@@
We gather we have a linear system $\left(\begin{matrix}c_{n-1}\\ \vdots\\ c_{1}\end{matrix}\right)=T^{-1}\left(\begin{matrix}f(1)\\ \vdots\\ f({n-1})\end{matrix}\right)$ where the matrix $T^{-1}$ appeared in the proof of Lemma 4.10. Abusing notation and putting $t^{R}_{k}:\mathbb{R}\to\mathbb{R}$ and $t^{L}_{k}:\mathbb{R}\to\mathbb{R}$ as in the proof of the same lemma, solving this system for the values of $f$ we obtain
$\left(\begin{matrix}f(1)\\ \vdots\\ f({n-1})\end{matrix}\right)$ $=T\left(\begin{matrix}c_{n-1}\\ \vdots\\ c_{1}\end{matrix}\right)=\sum_{k}c_{n-k}\left(\begin{matrix}t^{R}_{k}(n-1)\\ \vdots\\ t^{R}_{k}(1)\end{matrix}\right)=\sum_{k}c_{n-k}\left(\begin{matrix}t^{L}_{k}(1)\\ \vdots\\ t^{L}_{k}(n-1)\end{matrix}\right)$
$=\sum_{k}c_{n-k}\left(\begin{matrix}t^{R}_{n-k}(1)\\ \vdots\\ t^{R}_{n-k}(n-1)\end{matrix}\right)=\sum_{k}c_{k}\left(\begin{matrix}t^{R}_{k}(1)\\ \vdots\\ t^{R}_{k}(n-1)\end{matrix}\right)$
@@
There are also quantum versions of this duality. Since the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$ is a quantization of the coordinate ring $\mathcal{O}(\mathcal{R}_{\mathrm{SL}(n)})$ of the character variety, Fock–Goncharov–Shen duality suggests that the positive tropical integer points $\mathcal{A}^{+}_{\mathrm{PGL}(n)}(\mathbb{Z}^{t})$ should index the elements of a canonical basis of the skein algebra, in particular a canonical topological basis of $\mathcal{O}(\mathcal{R}_{\mathrm{SL}(n)})$ when $q^{1/n}=1$. Indeed, following Goncharov–Shen’s lead, upon choosing an ideal triangulation $\mathfrak{L}$ of $\mathfrak{S}$ we can think of the corresponding balanced Knutson–Tao fan $\mathcal{C}(\mathfrak{S},\mathfrak{L})$ as a model for $\mathcal{A}^{+}_{\mathrm{PGL}(n)}(\mathbb{Z}^{t})$. When the surface $\mathfrak{S}=\mathfrak{A}$ is the annulus with triangulation $\mathfrak{L}$ as above, then this model for $\mathcal{A}^{+}_{\mathrm{PGL}(n)}(\mathbb{Z}^{t})$, rather, the subset constrained to be zero on the boundary, is precisely the balanced Knutson–Tao fan $\mathcal{C}(\mathfrak{A},\mathfrak{L})$ appearing in Corollary 6.2, whose points index the monomial basis elements of the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{A})$ appearing in Theorem 2.1 via their highest degrees with respect to the quantum trace map $\tau^{q}_{\mathfrak{L}}$. The fact that these points appear as the highest polynomial degrees of the basis elements, as detected by the quantum trace map, is another indication of the tropical geometric nature of this duality.

### v2

@@ 4.4.5. Quantum trace map: definition, part 2 of 2
for $(a,b,c)$ interior, while for boundary vertices
$P(j(n-j)0,(j+1)(n-j-1)0)=P((n-j)0j,(n-j-1)0(j+1)=P(0j(n-j),0(j+1)(n-j-1))=+1$
@@
For $j=1,\dots,n-1$ and symbols $Z$ and $X$, define $n\times n$ diagonal and block diagonal matrices
$S^{\mathrm{edge}}_{j}(Z)$ $=Z^{-j/n}\mathrm{diag}\left(\begin{smallmatrix}Z&Z&\ldots&Z&1&1&\ldots&1\end{smallmatrix}\right),$
$S^{\mathrm{left}}_{j}(X)$ $=X^{-(j-1)/n}\mathrm{diag}\left(\begin{smallmatrix}X&\ldots&X&\left(\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right)&1&\ldots&1\end{smallmatrix}\right),$
$S^{\mathrm{right}}_{j}(X)$ $=X^{+(j-1)/n}\mathrm{diag}\left(\begin{smallmatrix}1&\ldots&1&\left(\begin{smallmatrix}1&0\\ 1&1\end{smallmatrix}\right)&X^{-1}&\ldots&X^{-1}\end{smallmatrix}\right)$
@@
$L^{\omega}_{i,j}$ $=\left[M^{\mathrm{edge}}(Z_{1},\dots,Z_{n-1})M^{\mathrm{left}}((X_{abc})_{abc})M^{\mathrm{edge}}(Z^{\prime}_{1},\dots,Z^{\prime}_{n-1})\right]_{\overline{i},\overline{j}},$
@@
Remark 4.8.
The definition of the quantum trace map discussed in this and the previous sub-sub-section reproduces, essentially on the nose, the $n=n$ construction of [50]; see also [13, 14].
@@ 4.5.2. Annulus: a piecewise linear computation
Let $L^{\omega}$ and $R^{\omega}$ be the standard quantum left and right matrices. In general, an entry $L^{\omega}_{ij}$ or $R^{\omega}_{ij}$ is a sum of monomials in $\mathcal{T}_{n}^{\omega}(\mathfrak{D}_{3})$. Put $M=L^{\omega}$, which recall is lower triangular. Note that similar facts hold for the right matrix as well, which we will not use.
@@
    $\mathrm{deg}\left(\prod_{i=n+1-k}^{n}M_{ii}\right)(a,b,c)$ $=\min\{f_{k}(n-a),g_{k}(n-a)\},$
@@
(1) By definition, $M$ is the product of lower triangular, and diagonal, matrices $M_{v}$, $v\in\Theta_{n}^{\prime}$, each of whose $ij$-entry is of the form $X_{v}^{d_{vij}/n}$ for some $d_{vij}\in\mathbb{Z}$, as well as lower triangular matrices involving no variables. Rather, $M$ is the Weyl ordering of such a product, but the Weyl ordering will play no role throughout this proof since we are only interested in degrees, so we will omit it for notational simplicity.
(2) This is one of the relations appearing in the main result of [13].
(3) The first statement is immediate from the form of the matrices $M_{v}$. The second statement is because the off diagonal terms $X_{v}^{d_{vij}/n}$ of the $M_{v}$ have exponent $d_{vij}<0$.
(4) Clearly $M$ does not involve the $X_{0j(n-j)}$, so assume $a\geq 1$. Recall that the $(\overline{m},\overline{m})$ entry of $M$ is the $(m,m)$ entry of the product of matrices $S^{\mathrm{edge}}_{l}(X_{l0(n-l)})$ and $S^{\mathrm{edge}}_{l}(X_{l(n-l)0}$ for $11\leq l\leq n-$ 1 and $S^{\mathrm{left}}_{j}(X_{(j-1)(n-i)(i-j+1)})$ for $12\leq j\leq i\leq n-$ 1, and $S^{\mathrm{left}}_{1}$ a number of times. We only need to observe then that, in the former, $X_{l0(n-l)}$ and $X_{l(n-l)0}$ appear with power $1-\ell/n$ (resp. $-\ell/n$) when $m\leq l$ (resp. $m>l$), and, in the latter, $X_{(j-1)(n-i)(i-j+1)}$ appears with power $1-(j-1)/n$ (resp. $-(j-1)/n$) when $m\leq j-1$ (resp. $m>j-1$).
(5) This is part of the mathematical content of [50, Equation (170) and Theorem 10.2]; for an earlier appearance of these piecewise linear formulas, see [71, Theorem 8.22]. For the sake of self-containedness, we provide this short calculation here as well. The formulas work when $a=0$, so assume $a\geq 1$. Note that the larger degree $n-a$ for $M_{ii}$ occurs if and only if $n+1-a\leq i$, else it has the smaller degree $-a$.
@@
Remark 4.11.
It has been known by experts for some time now, see for example the Acknowledgements and the comment about [71] in the proof of item (5) above as well as the upcoming work [48], that a certain set of equivalence classes of $n$-valent graphs in a punctured surface $\mathfrak{S}$ equipped with an ideal triangulation $\mathfrak{L}$ is in natural one-to-one correspondence with a certain subset of $\mathbb{N}^{N}$ where $N$ depends only on $n$ and the topology of $\mathfrak{S}$; see also [77]. Here, the graph is called a ‘reduced web’ and the corresponding $N$-tuple constitutes its ‘tropical web coordinates’. Such a result generalizes to the case $n=n$ the work [19, 20] in the case $n=2$ and the work [16, 17, 24] in the case $n=3$.
When the surface $\mathfrak{S}=\mathfrak{A}$ is the annulus, the reduced webs represent the same skein basis webs being studied in the present paper. Although, technically, we do not explicitly use this $n=n$ result about tropical web coordinates, it essentially underlies our entire strategy. This is because the $N$-tuple of tropical web coordinates is precisely the $N$-tuple of exponents of the highest term of the quantum trace polynomial of the basis web, which we compute in the current section. See [5, 39] for the corresponding quantum trace calculations for $n=2$ and $n=3$ for any surface $\mathfrak{S}$.
Lastly, building on what was learned from the $n=2$ and $n=3$ cases, we owe the majority of our understanding of Fock–Goncharov duality in the $n=n$ case, discussed in Section 6, to conversations with Zhe Sun; see the Acknowledgements.
@@ 4.5.3. Annulus: finishing
Abusing notation and putting $t^{R}_{k}:\mathbb{R}\to\mathbb{R}$ defined by $t^{R}_{k}(a)=\min\{f_{k}(n-a),g_{k}(n-a)\}$, it suffices to show that the $n-1$ vectors $(t^{R}_{k}(n-1),\dots,t^{R}_{k}(2),t^{R}_{k}(1))$ in $\mathbb{R}^{n-1}$ varying over $k=1,2,\dots,n-1$ are linearly independent. Indeed, if we arrange these $n-1$ vectors as the columns of a $(n-1)\times(n-1)$ matrix $T^{R}$, $T^{R}_{ij}=t^{R}_{j}(n-i)$, then this matrix is the inverse of the matrix $\left(T^{R}\right)^{-1}=(1/n)\left(\begin{smallmatrix}&&&&0&0&-1&2\\ &&&&0&-1&2&-1\\ &&&&-1&2&-1&0\\ \vdots&\vdots&\vdots&\reflectbox{$\ddots$}&&&&\\ -1&2&-1&\ldots&&&&\\ 2&-1&0&\ldots&&&&\end{smallmatrix}\right)$. ∎
@@ 5.1. Embedding the commutative polynomial algebra into quantum tori
In other words, for every such $q^{1/n}$ we obtain an explicit geometric embedding of the polynomial algebra $\mathcal{S}_{n}^{q}\cong\mathbb{C}[x_{1},x_{2},\dots,x_{n-1}]$ into the quantum torus $\mathcal{T}_{n}^{\omega}(\mathfrak{D}_{3}^{L})\otimes\mathcal{T}_{n}^{\omega}(\mathfrak{D}_{3}^{R})$.
@@
By [69, Corollary 20], see also [6, 62] for the case $n=2$, the skein module $\mathcal{S}_{n}^{q}(\mathfrak{M})$ of a three-dimensional manifold $\mathfrak{M}$ (Remark 1.5) evaluated at $q^{1/n}=1$ is isomorphic to the, a priori, unreduced coordinate ring over $\mathbb{C}$ of the $\mathrm{SL}(n)$ character variety of $\mathfrak{M}$. Said another way, the reduced coordinate ring of the $\mathrm{SL}(n)$ character variety of $\mathfrak{M}$ is isomorphic to $\mathcal{S}_{n}^{q}(\mathfrak{M})(q^{1/n}=1)/\sqrt{0}$, namely the quotient of $\mathcal{S}_{n}^{q}(\mathfrak{M})(q^{1/n}=1)$ by its set of nilpotent elements. By Corollary 5.3, since $q^{1/n}=1\notin Q$, taking $\mathfrak{M}=\mathfrak{A}\times\mathfrak{I}$, the coordinate ring of the $\mathrm{SL}(n)$ character variety of the annulus is reduced and isomorphic to the skein algebra $\mathcal{S}_{n}^{q}(q^{1/n}=1)\cong\mathbb{C}[x_{1},x_{2},\dots,x_{n-1}]$.
@@ 5.2. Injectivity of the splitting map for the twice punctured sphere
The argument is identical to that of the $n=3$ result [35, Theorem 8.1], or the earlier $n=2$ result of [47], both proved for any surface $\mathfrak{S}$, once it is known that the monomial webs $(B^{\circ}_{1})^{\prime m_{1}}(B^{\circ}_{2})^{\prime m_{2}}\cdots(B^{\circ}_{n-1})^{\prime m_{n-1}}$ in the stated skein algebra $\widetilde{\mathcal{S}}_{n}^{q}(\overline{\mathfrak{S}})$, which are the images under the splitting map of the monomial basis webs $(B^{\circ}_{1})^{m_{1}}(B^{\circ}_{2})^{m_{2}}\cdots(B^{\circ}_{n-1})^{m_{n-1}}$ in $\left(\mathcal{S}_{n}^{q}\right)^{\circ}$, are linearly independent in $\widetilde{\mathcal{S}}_{n}^{q}(\overline{\mathfrak{S}})$. This was essentially proved above as part of the proof of the main theorem. Technically, this independence was proved for the annular surface $\overline{\overline{\mathfrak{S}}}$ obtained from $\overline{\mathfrak{S}}$ by cutting out another monoangle from the remaining internal puncture, but since the monomial webs in the surface $\overline{\overline{\mathfrak{S}}}$ are the images under the splitting map of the monomial webs in the surface $\overline{\mathfrak{S}}$, the latter are independent as well. Note also that, when mimicking the argument of [35], the $n=3$ ‘cut’ and ‘glue’ maps of [35, Section 7] are replaced by the corresponding $n=n$ maps in [49, Section 7.4]. ∎
@@ 6. Fock–Goncharov duality
The balanced Knutson–Tao lattice (resp. fan) is the set of functions $f:\Theta_{n}^{\prime}\to\mathbb{R}$ satisfying the property that all the rhombus numbers $r(f)^{\mathrm{t}}_{ij}$ and $r(f)^{\mathrm{bl}}_{ij}$ and $r(f)^{\mathrm{br}}_{ij}$ are integers (resp. natural numbers) [43]. Note that the balanced Knutson–Tao lattice (resp. fan) is a subgroup (resp. sub-monoid) of $\mathbb{R}^{\Theta_{n}^{\prime}}$ under addition, in particular it contains the zero function. We denote by $\mathcal{C}$ the sub-monoid that is the intersection of the balanced Knutson–Tao fan with those functions $f$ satisfying $f(a,b,c)=0$ for all $a=0$.
Let the functions $t^{L}_{k}$ and $t^{R}_{k}$ in $\mathbb{N}^{\Theta_{n}^{\prime}}$ for $k=1,2,\dots,n-1$ be as in Section 4.5.3. One checks $t^{R}_{k}\in\mathcal{C}$ for each $k$. By the symmetry $t^{L}_{n-k}=t^{R}_{k}$ this is also true for the $t^{L}_{k}$. Moreover, we have the following specific values for the rhombus numbers: $r(t^{R}_{k})^{\mathrm{bl}}_{kj}=1$ for all $j=1,2,\dots,n-k$ and all other rhombus numbers $r(t^{R}_{k})^{\mathrm{bl}}_{ij}$, $i\neq k$, and $r(t^{R}_{k})^{\mathrm{t}}_{ij}$ and $r(t^{R}_{k})^{\mathrm{br}}_{ij}$ are zero.
@@
We gather we have a linear system $\left(\begin{smallmatrix}c_{n-1}\\ \vdots\\ c_{1}\end{smallmatrix}\right)=T^{-1}\left(\begin{smallmatrix}f(1)\\ \vdots\\ f({n-1})\end{smallmatrix}\right)$ where the matrix $T^{-1}$ appeared in the proof of Lemma 4.12. Abusing notation and putting $t^{R}_{k}:\mathbb{R}\to\mathbb{R}$ and $t^{L}_{k}:\mathbb{R}\to\mathbb{R}$ as in the proof of the same lemma, solving this system for the values of $f$,
$\left(\begin{smallmatrix}f(1)\\ \vdots\\ f({n-1})\end{smallmatrix}\right)$ $=T\left(\begin{smallmatrix}c_{n-1}\\ \vdots\\ c_{1}\end{smallmatrix}\right)=\sum_{k}c_{n-k}\left(\begin{smallmatrix}t^{R}_{k}(n-1)\\ \vdots\\ t^{R}_{k}(1)\end{smallmatrix}\right)=\sum_{k}c_{n-k}\left(\begin{smallmatrix}t^{L}_{k}(1)\\ \vdots\\ t^{L}_{k}(n-1)\end{smallmatrix}\right)$
$=\sum_{k}c_{n-k}\left(\begin{smallmatrix}t^{R}_{n-k}(1)\\ \vdots\\ t^{R}_{n-k}(n-1)\end{smallmatrix}\right)=\sum_{k}c_{k}\left(\begin{smallmatrix}t^{R}_{k}(1)\\ \vdots\\ t^{R}_{k}(n-1)\end{smallmatrix}\right)$
@@
There are also quantum versions of this duality. Since the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{S})$ is a quantization of the coordinate ring $\mathcal{O}(\mathcal{R}_{\mathrm{SL}(n)})$ of the character variety, Fock–Goncharov–Shen duality suggests that the positive tropical integer points $\mathcal{A}^{+}_{\mathrm{PGL}(n)}(\mathbb{Z}^{t})$ should index the elements of a canonical basis of the skein algebra, in particular a canonical topological basis of $\mathcal{O}(\mathcal{R}_{\mathrm{SL}(n)})$ when $q^{1/n}=1$. Indeed, following Goncharov–Shen’s lead, upon choosing an ideal triangulation $\mathfrak{L}$ of $\mathfrak{S}$ we can think of the corresponding balanced Knutson–Tao fan $\mathcal{C}(\mathfrak{S},\mathfrak{L})$ as a model for $\mathcal{A}^{+}_{\mathrm{PGL}(n)}(\mathbb{Z}^{t})$. When the surface $\mathfrak{S}=\mathfrak{A}$ is the annulus with triangulation $\mathfrak{L}$ as above, then this model for $\mathcal{A}^{+}_{\mathrm{PGL}(n)}(\mathbb{Z}^{t})$, rather, the subset constrained to be zero on the boundary, is precisely the balanced Knutson–Tao fan $\mathcal{C}(\mathfrak{A},\mathfrak{L})$ appearing in Corollary 6.2, whose points index the monomial basis elements of the skein algebra $\mathcal{S}_{n}^{q}(\mathfrak{A})$ appearing in Theorem 2.1 via their highest degrees with respect to the quantum trace map $\tau^{q}_{\mathfrak{L}}$. That these points appear as the highest polynomial degrees of the basis elements, as detected by the quantum trace map, is another indication of the tropical geometric nature of this duality.